OpenStudy (anonymous):

if a wall is measured to 12.2 meters and on a drawing it is measured to 3.25in, what is a wall that is 5.50in in meters?

OpenStudy (anonymous):

3.25 / 5.5 = 12.2 / x x = 5.5 * 12.2 / 3.25

OpenStudy (anonymous):

12.2 meters are represented by 3.25 inches. So 12.2/3.25 is how many meters would ONE inch represent. Now to get how many meters are represented in 5.5 inches, just multiply times 5.5 Hope that makes it easier to grasp.

OpenStudy (anonymous):

x = 20.65